In early version of double slit experiment discussed,carnal and mlynek use helium atoms ( from reservoir maintained at 295k) that exit the nozzle after expansion with wavelength = 0.56 A (a) what is the speed of the helium atoms


1
Expert's answer
2021-07-19T09:50:58-0400

"E_1=E_2+E_3,"

"\frac{mv^2}{2}=\frac{hc}{\lambda}+\frac{3kT}2,"

"v=\sqrt{\frac{2hc+3k\lambda T}{m\lambda}},"

"v=1.034\cdot 10^6~\frac ms."